# Trustworthy Transfer Learning: A Survey

### v1

@@ 6.1 Benchmarking Negative Transfer
$\mathcal{E}_{T}\left(A^{\mathrm{tl}}\left(D_{S},D_{T}\right)\right)>\mathcal{E}_{T}\left(A^{\mathrm{tl}}\left(\emptyset,D_{T}\right)\right)$ (25)
@@
This definition reveals (?, ?, ?) that given a learning algorithm, there are two major factors determining if negative transfer occurs: the distribution discrepancy between source and target domains and the size of the labeled target data. Negative transfer has been observed theoretically and empirically in various applications (?, ?, ?, ?). Recent work (?) also explores identifying and characterizing the failure modes that pre-training can and cannot address. Despite the extensive work on transfer learning techniques, up until now, little effort (if any) has been devoted to rigorously understanding the boundary between positive and negative transfer given a learning algorithm. It remains an open question to determine when the negative transfer will occur given finite source and target samples (or a source model and finite target samples). Therefore, rather than focusing on performance improvement, more efforts can be dedicated to benchmarking the negative transfer of transfer learning models, e.g., the change from positive to negative transfer can be affected by the magnitude of distribution shifts and the number of target samples. This could provide valuable insights into when a transfer learning model can work well for real-world applications.
6.2 Cross-modal Transferability
@@
Cross-modal transfer learning (?, ?, ?) aims at understanding knowledge transferability when the source and target domains have different types of data modalities, e.g., transferring knowledge from a text-based source domain to an image-based target domain. This differs from multi-modal learning (?, ?) which maps different data modalities in a unified latent feature space over pair-wise training samples. In contrast, cross-modal transfer learning focuses on investigating what knowledge can be transferred across data modalities. Although large language models (LLMs) have been applied to various scientific discovery tasks (?), it is unclear what knowledge is being transferred in this process. Additionally, there is a lack of theoretical understanding regarding how LLMs generalize to downstream tasks with different data modalities.
6.3 Physics-Informed Transfer Learning
@@
Physics-informed machine learning (?, ?) aims to improve the training of machine learning models by incorporating physical domain knowledge as soft constraints on an empirical loss function. This alleviates the need for a large amount of high-quality data when training deep neural networks to solve scientific problems. Recent studies have introduced transfer learning to understand the knowledge transferability of physics-informed neural networks across tasks. For example, ? (?) study the transferability of physics-informed neural networks across differential equations. ? (?) and ? (?) study the transfer learning performance of deep operator networks across partial differential equations. ? (?) further analyze the transfer behavior of neural operators pre-trained on a mixture of different physics problems. However, the theoretical explanation regarding the generalization performance of physics-informed neural networks under distribution shifts is unclear. Besides, it can be seen that in the context of transfer learning, the source knowledge can be provided from multiple aspects, including labeled source samples (?) (e.g., Subsection 3.1.1), pre-trained source models (?) (e.g., Subsection 4.1.1), synthetic data generated by physics-based simulators (?) (e.g., Subsection 5.5), and fundamental physical rules (?). This can motivate a generic physics-informed transfer learning problem involving multi-faceted knowledge transfer from the source to the target domains.
6.4 Trade-off between Transferability and Trustworthiness
@@
In standard machine learning, the trade-off between prediction accuracy and trustworthiness has been theoretically studied e.g., accuracy vs. group fairness (?, ?), accuracy vs. adversarial robustness (?, ?, ?), accuracy vs. privacy (?), accuracy vs. explainability (?), etc. It has been observed that trustworthiness may be at odds with the prediction accuracy in a single domain. In contrast, recent works (?, ?) reveal that both trustworthiness (e.g., adversarial robustness) and prediction accuracy can be improved in the target domain by leveraging relevant knowledge from source domains. This motivates us to re-think the fundamental trade-off between knowledge transferability and trustworthiness in the context of transfer learning. Specifically, there are several open questions: (1) Can source knowledge consistently enhance trustworthiness and transfer accuracy in the target domain under various distribution shifts and data modalities? (2) Is there an inherent trade-off between trustworthiness and transfer accuracy in the target domain when discovering and transferring knowledge from the source data/model? These studies will significantly expand the application of transfer learning techniques by clarifying when trained models can be trusted and how well they perform.
7 Conclusion

### v2

@@ 6.1. Benchmarking Negative Transfer
(25) $\mathcal{E}_{T}\left(A^{\mathrm{tl}}\left(D_{S},D_{T}\right)\right)>\mathcal{E}_{T}\left(A^{\mathrm{tl}}\left(\emptyset,D_{T}\right)\right)$
@@
This definition reveals (A_distance; KuzborskijO13; WangDPC19) that given a learning algorithm, there are two major factors determining if negative transfer occurs: the distribution discrepancy between source and target domains and the size of the labeled target data. Negative transfer has been observed theoretically and empirically in various applications (rosenstein2005transfer; Ben-DavidLLP10; 0002CZG19; WangDPC19). Recent work (cohen2024ask) also explores identifying and characterizing the failure modes that pre-training can and cannot address. Despite the extensive work on transfer learning techniques, up until now, little effort (if any) has been devoted to rigorously understanding the boundary between positive and negative transfer given a learning algorithm. It remains an open question to determine when the negative transfer will occur given finite source and target samples (or a source model and finite target samples). Therefore, rather than focusing on performance improvement, more efforts can be dedicated to benchmarking the negative transfer of transfer learning models, e.g., the change from positive to negative transfer can be affected by the magnitude of distribution shifts and the number of target samples. This could provide valuable insights into when a transfer learning model can work well for real-world applications.
6.2. Cross-modal Transferability
@@
Cross-modal transfer learning (ShenLDSKNT23; DinhZZLGRSP022; socher2013zero) aims at understanding knowledge transferability when the source and target domains have different types of data modalities, e.g., transferring knowledge from a text-based source domain to an image-based target domain. This differs from multi-modal learning (RadfordKHRGASAM21; huang2021makes), which maps different data modalities in a unified latent feature space over pair-wise training samples. In contrast, cross-modal transfer learning focuses on investigating what knowledge can be transferred across data modalities. Although large language models (LLMs) have been applied to various scientific discovery tasks (DinhZZLGRSP022), it is unclear what knowledge is being transferred in this process. Additionally, there is a lack of theoretical understanding regarding how LLMs generalize to downstream tasks with different data modalities.
6.3. Physics-informed Transfer Learning
@@
Physics-informed machine learning (karniadakis2021physics; raissi2019physics) aims to improve the training of machine learning models by incorporating physical domain knowledge as soft constraints on an empirical loss function. This alleviates the need for a large amount of high-quality data when training deep neural networks to solve scientific problems. Recent studies have introduced transfer learning to understand the knowledge transferability of physics-informed neural networks across tasks. For example, (desai2022oneshot) study the transferability of physics-informed neural networks across differential equations. (goswami2022deep; xu2023transfer) study the transfer learning performance of deep operator networks across partial differential equations. (subramanian2023towards) further analyze the transfer behavior of neural operators pre-trained on a mixture of different physics problems. However, the theoretical explanation regarding the generalization performance of physics-informed neural networks under distribution shifts is unclear. Besides, it can be seen that in the context of transfer learning, the source knowledge can be provided from multiple aspects, including labeled source samples (wiles2022a) (e.g., Subsection 3.1.1), pre-trained source models (GPT4) (e.g., Subsection 4.1.1), synthetic data generated by physics-based simulators (andrychowicz2020learning) (e.g., Subsection 5.5), and fundamental physical rules (karniadakis2021physics). This can motivate a generic physics-informed transfer learning problem involving multi-faceted knowledge transfer from the source to the target domains.
6.4. Trade-off between Transferability and Trustworthiness
@@
In standard machine learning, the trade-off between prediction accuracy and trustworthiness has been theoretically studied, e.g., accuracy vs. group fairness (ZhaoG19; DuttaWYC0V20), accuracy vs. adversarial robustness (TsiprasSETM19; ZhangYJXGJ19; YangRZSC20), accuracy vs. privacy (BiettiWD0W22), accuracy vs. explainability (ZarlengaBCMGDSP22), etc. It has been observed that trustworthiness may be at odds with the prediction accuracy in a single domain. In contrast, recent work (SalmanIEKM20; davchev2019empirical) reveals that both trustworthiness (e.g., adversarial robustness) and prediction accuracy can be improved in the target domain by leveraging relevant knowledge from source domains. This motivates us to rethink the fundamental trade-off between knowledge transferability and trustworthiness in the context of transfer learning. Specifically, there are several open questions: (1) Can source knowledge consistently enhance trustworthiness and transfer accuracy in the target domain under various distribution shifts and data modalities? (2) Is there an inherent trade-off between trustworthiness and transfer accuracy in the target domain when discovering and transferring knowledge from the source data/model? These studies will significantly expand the application of transfer learning techniques by clarifying when trained models can be trusted and how well they perform.
6.5. Trustworthy Transfer Learning of Foundation Models
Although foundation models have achieved surprising performance across high-impact domains, the underlying properties governing their behavior are not yet fully understood (bengio2024managing; kapoor2024position), especially when adapting them to downstream tasks. In this work, we summarize the key open questions from the perspective of trustworthy transfer learning. On one hand, knowledge transferability involves whether foundation models can be successfully adapted to downstream tasks with guaranteed performance. This post-training process leads to the following open questions:
  • Scaling Behaviors of Fine-tuning: Recent work (tay2022scale; zhang2024when) has investigated the power-law scaling relationship between LLM fine-tuning performance and model/data size. In contrast, (ren2025learning) provide a step-wise analysis of the learning dynamics for LLM fine-tuning under gradient descent. However, theoretical understanding still lags behind empirical findings, e.g., how generalization performance of fine-tuning is bounded in terms of model architecture, model and data scale, how learning dynamics vary across different fine-tuning (or other post-training) strategies, and how learning dynamics of fine-tuning explain the potential negative transfer behaviors.
  • Transferability Metrics for Foundation Model Selection: Recent work (lin2024selecting) leverages rectified scaling laws to predict fine-tuning performance, thereby facilitating model selection for LLMs. However, it is computationally expensive due to the need for fine-tuning in low-data regimes, and the theoretical foundations connecting scaling laws to fine-tuning performance are still underexplored. To enable model selection and customization in the AI model market (pei2023data), there is a growing need for theory-grounded and computationally efficient transferability metrics. In particular, it is feasible and desirable to establish principled guidelines for selecting optimal foundation models for specific target tasks, as well as identifying the application domains in which a given foundation model is best suited for fine-tuning.
  • AI Model Collapse: Generally, model collapse is defined as “a degenerative process affecting generations of learned generative models, in which the data they generate end up polluting the training set of the next generation” (shumailov2024ai). This phenomenon motivates us to rethink the impact of synthetic data (maybe indistinguishable from human-generated content) on both the pre-training and fine-tuning of foundation models. From the perspective of transfer learning, a key challenge for mitigating model collapse lies in handling open-world distribution shifts between real and AI-generated data.
On the other hand, the knowledge trustworthiness of foundation models in the transfer learning process might involve the following open questions:
  • Data Poisoning and Model Stealing Attacks: It is shown (WanWSK23; bowen2024scaling) that larger foundation models are significantly more susceptible to data poisoning attacks than smaller ones during fine-tuning. However, limited efforts have been devoted to theoretically understanding the connections between model size and adversarial vulnerability of foundation models in the context of transfer learning. In addition, recent work (carlini2024stealing) discusses the model stealing attacks to extract precise information from black-box production language models via APIs. These emerging threats highlight significant security challenges for the development and deployment of foundation models in both academia and industry. As a result, advancing certified robust techniques is essential to improve the reliability and trustworthiness of foundation models.
  • Transparency and Mechanistic Interpretability. Mechanistic interpretability aims at understanding the internal mechanisms of foundation models (dunefsky2024transcoders; ferrando2025do), thereby providing insights into their decision-making processes and enhancing model transparency. Nevertheless, due to the architectural complexity and scale of foundation models, it is challenging to systematically explain their emergent behaviors and key mechanisms behind their advanced capabilities, when they are adapted to open-world environments. Although fragmented efforts have been made, a systematic framework is still needed to advance the mechanistic interpretability of foundation models.
  • Ethical and Societal Impact in Foundation Model Customization: The ethical and societal implications, such as misinformation and privacy concerns, of open foundation models have been examined from the perspectives of AI developers (nikolinakos2023ethical; kapoor2024position). As these models are increasingly fine-tuned and customized by millions of non-expert users, a critical gap emerges between expert understanding and everyday use. This gap highlights the urgent need for intuitive risk assessments and protective measures that empower non-AI users to identify and mitigate intentional or unintentional misuse of foundation models.
7. Conclusion
@@
Acknowledgements.
This work is supported by National Science Foundation under Award No. IIS-2117902, and Agriculture and Food Research Initiative (AFRI) grant no. 2020-67021-32799/project accession no.1024178 from the USDA National Institute of Food and Agriculture. The views and conclusions are those of the authors and should not be interpreted as representing the official policies of the funding agencies or the government.